OpenStudy (anonymous):

6g^2 +17gd +12d^2 factor completely and simplify

OpenStudy (anonymous):

(2g+3d)(3g+4d) factored

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

:) thats simplyed already

myininaya (myininaya):

6g^2+17gd+12d^2 6g^2+17dg+12d^2 Ag^2+Bg+C so A=6, B=17d, C=12d^2 The trick is to find two factors of A*C that multiply to be A*C and add up to be B. A*C=(6)*(12d^2)=6d*12d=18d*4d=9d*8d and B=17d=8d+9d so replace Bg=17dg with 8dg+9dg so we have 6g^2+8dg+9dg+12d^2 now we factor by grouping (6g^2+8dg)+(9dg+12d^2) in the first paranthesis what do 6g^2 and 8dg have in common they have 2g in common so we can factor 2g out of first paranthesis so we have 2g(3g+4d)+(9dg+12d^2) now what do 9dg and 12d^2 have in common they have 3d in common so factor that out in the second paranthesis so we have 2g(3g+4d)+3d(3g+4d) so since we have 3g+4d in both terms we can factor that our (3g+4d)(2g+3d)